High buffer capacity: NaH2PO4-H2O (1.3804g) K2HPO4(0.551g)

Target pH=6.7

Volume of buffer=40mL

1. using activity coefficients to calculate the pH

2. calculate buffer capacity against acids(1mL 0.1M HCl)

3. calculate buffer capacity against bases(1mL 0.1M NaOH)
Answer #1

The no.of moles of NaH2PO4.H2O = 1.3804 g/(138 g/mol) ~ 10^(-2) mol

The no.of moles of K2HPO4 = 0.551 g/(174.2 g/mol) ~ 0.3×10^(-2) mol

According to Henderson-Hasselbulch equation:

pH = pKa + Log([K2HPO4]/[NaH2PO4.H2O])

= 7.2 + Log{0.3×10^(-2)/10^(-2)}

= 6.68

2. The no.of moles of added HCl = 0.01×10^(-2) mol

The resulting no.of moles of base = 0.3 - 0.01 = 0.29×10^(-2) mol

The resulting no.of moles of acid = 1 + 0.01 = 1.01×10^(-2) mol

Therefore, pH = 7.2 + Log(0.29/1.01) = 6.66

3. The no.of moles of added NaOH = 0.01×10^(-2) mol

The resulting no.of moles of base = 0.3 + 0.01 = 0.31×10^(-2) mol

The resulting no.of moles of acid = 1 - 0.01 = 0.99×10^(-2) mol

Therefore, pH = 7.2 + Log(0.31/0.99) = 6.70